OpenStudy (anonymous):

Air is being pumped into a balloon, at a rate represented by the equation y = t1/3 where y is the rate of airflow into the balloon in cm3/s and t is the time in seconds. The balloon already has 10 cm3 of air in it before air is pumped into it. The balloon will burst if there is more than 100cm3 of air in it. Find the time taken for the balloon to burst.

OpenStudy (anonymous):

more than 270 seconds

OpenStudy (anonymous):

can u plz explain

OpenStudy (anonymous):

i explained in the previous post you started for this thread.

OpenStudy (anonymous):

90 = t/3 t = 270 seconds

OpenStudy (anonymous):

there is already 10 cm^3 of air in the balloon, so 90cm^3 of air will be enough to make it burst so 90 = t/3 so for t > 270, the balloon will burst